OpenStudy (anonymous):

A triangular prism has a length of 6 inches. A larger similar triangular prism has a length of 20 inches. If 270 square inches of paper is used to wrap the smaller prism with minimum wastage, approximately how many square inches of paper will be required to wrap the larger prism? Answer 300 square inches 900 square inches 1000 square inches 3000 square inches

OpenStudy (jlastino):

This is a proportion 270/6 = x/20 Solve for x

OpenStudy (anonymous):

x=900 square inches

Directrix (directrix):

If two polyhera are similar, the square of their scale factor is equal to the ratio of their surface areas. (6/20) ^2 = 270/A where A is the surface area of the larger polydron. (3/10)^2 = 270/A 9/100 = 270/A A = 3000 square inches